OpenStudy (anonymous):

What is a solution of x^2+6x=-5? A) x= -6 B) x= -1 C) x= 1 D) x=6

OpenStudy (amistre64):

try and error ....

OpenStudy (anonymous):

im not exactly sure how.

OpenStudy (amistre64):

what is x^2 + 6x ... when x=-6? ... when x=-1? ... when x= 1? ... when x= 6?

OpenStudy (amistre64):

if you dont know how to add and multiply yet, you may want to review those operations

OpenStudy (mathmale):

Of course this direct substitution of the possible answers will "work." But it'd be more interesting to actually solve the quadratic equation x^2+6x=-5, or x^2+6x+5=0.

OpenStudy (anonymous):

i dont know how im supposed to solve this. with the variables.

OpenStudy (amistre64):

there are many methods that can be used, and your saying you are not aware of any of them. you should prolly review your material and ask questions about the processes that dont make sense.

OpenStudy (mathmale):

You have at least two choices: 1) Take each of the four possible answers and substitute each into the original equation, as amistre64 has suggested. If the equation is true, you've found a solution. 2) Actually solve the given quadratic equation. Please decide which approach YOU want to take.

OpenStudy (amistre64):

an example of trial and error: is x=-2 a solution to: 3x^2 - 2x = 7? try it, everywhere there is an x, replace it by -2: 3(-2)^2 - 2(-2) = 7 3(4) + 4 = 7 12 + 4 = 7 16 = 7 since 16 is not 7, -2 is not a solution.

OpenStudy (anonymous):

so if there are four variables would they all be replaced by a number?

OpenStudy (mathmale):

x=6 is one of the four possible solutions. Would you, @Book_Nerd_101, now please follow amistre64's example, above, substituting x=6 instead of his x=-2? Respond by saying "the equation is true" or "the equation is false". If the equation is true, x=6 is a solution. If the equation is false, x=6 is not a solution. Do this, please.

OpenStudy (amistre64):

they would be replaced by their tried values yes. if they are all an x, or some power of x, then you would replace them all by a tried value of x.

OpenStudy (anonymous):

ok. it makes more sense now. thank you.

OpenStudy (amistre64):

if you work out the solution, we can dbl chk your process :) good luck

OpenStudy (anonymous):

Tips to quickly solve these types of quadratic equations ax^2 + bx + c = 0 TIP 1. When a + b + c = 0, one real root is (1) and the other is c/a. Example; 7x^2 -13x + 6 = 0 Since a + b + c = 0, then one real root is (1) and the other is (c/a = 6/7). TIP 2. When a - b + c = 0, one real root is (-1) and the other is (-c/a) Above example: x^+ 6x + 5 = 0. Since a - b + c = 0, one real root is (-1) and the other is (-c/a = -5) Memorize these TIPS, They will save you a lot of time.